# GaussianFormer3D: Multi-Modal Gaussian-based Semantic Occupancy Prediction with 3D Deformable Attention

### v1

@@ Abstract
3D semantic occupancy prediction is critical for achieving safe and reliable autonomous driving. Compared to camera-only perception systems, multi-modal pipelines, especially LiDAR-camera fusion methods, can produce more accurate and detailed predictions. Although most existing works utilize a dense grid-based representation, in which the entire 3D space is uniformly divided into discrete voxels, the emergence of 3D Gaussians provides a compact and continuous object-centric representation. In this work, we propose a multi-modal Gaussian-based semantic occupancy prediction framework utilizing 3D deformable attention, named as GaussianFormer3D. We introduce a voxel-to-Gaussian initialization strategy to provide 3D Gaussians with geometry priors from LiDAR data, and design a LiDAR-guided 3D deformable attention mechanism for refining 3D Gaussians with LiDAR-camera fusion features in a lifted 3D space. We conducted extensive experiments on both on-road and off-road datasets, demonstrating that our GaussianFormer3D achieves high prediction accuracy that is comparable to state-of-the-art multi-modal fusion-based methods with reduced memory consumption and improved efficiency. Project website: GaussianFormer3D.
footnotetext: Corresponding author.
1 Introduction
@@
Perception systems are essential for constructing safe and reliable autonomous vehicles [19, 89]. Among various perception tasks, 3D semantic occupancy prediction [5, 24, 22, 88, 70, 36] is particularly crucial as it enables fine-grained understanding of both semantics and geometry information of the environments. Recent advances in vision-based occupancy prediction have demonstrated impressive performance on large-scale datasets [4, 1, 39, 62]. However, camera sensors are sensitive to lighting conditions and lack accurate depth estimation, motivating researchers to incorporate other sensor modalities to enhance the robustness of autonomous driving perception systems.
Figure 1: We propose a LiDAR-camera fusion-based semantic occupancy prediction framework named GaussianFormer3D. We use 3D Gaussians instead of dense grids to reduce memory consumption and enhance algorithm efficiency. GaussianFormer3D achieves comparable performance to state-of-the-art multi-modal occupancy methods with reduced memory usage.
@@
LiDAR sensors have been widely applied to autonomous driving for perception tasks such as 3D object detection [93, 77, 81, 59, 8, 30]. Compared to cameras, LiDAR provides more accurate depth information and captures finer geometric relationships of objects, making LiDAR particularly advantageous for 3D semantic occupancy prediction [57, 58, 76, 10, 80, 73, 95, 46, 6, 75, 71]. However, LiDAR-based pipelines may struggle to capture accurate semantic information for small objects, where camera-based methods excel. To balance semantics and geometry, multi-modal fusion-based algorithms have been proposed to leverage the strengths of different sensors. Fusion-based semantic occupancy prediction methods include LiDAR-camera fusion [52, 66, 32, 3, 63, 60, 87], camera-radar fusion [72, 45] and LiDAR-camera-radar fusion [51]. Among these sensor configurations, LiDAR-camera fusion is the most popular and top-performing one.
@@
Most existing LiDAR-camera occupancy networks employ a 3D voxel-based [52, 66, 32, 87] or a 2D bird’s-eye-view (BEV)-based representation [60], both depicting a 3D scene as a dense grid-based structure. Despite achieving comparable performance, they inevitably struggle with redundant empty grids and high computational costs. Recently, inspired by the success of 3D Gaussian splatting [28], a novel object-centric representation has been explored for the first time in vision-based 3D semantic occupancy prediction [25, 23, 13]. GaussianFormer [25, 23] represents a 3D scene as a set of 3D Gaussians, each consisting of a mean, covariance and semantic label. These Gaussians are refined with a 2D deformable attention mechanism [94], and then processed by an efficient Gaussian-to-voxel splatting module to predict the semantic occupancy. However, current Gaussian-based methods [25, 23] rely solely on 2D image feedback to update 3D Gaussians, limiting their ability to model 3D space with accurate depth information and fine-grained geometric structure. How to exploit data from other modalities, such as LiDAR, to refine and obtain a more accurate 3D Gaussian representation for efficient semantic occupancy prediction remains unexplored.
@@
Based on the aforementioned observations, we propose GaussianFormer3D: a multi-modal Gaussian-based semantic occupancy prediction framework with 3D deformable attention, as shown in Fig. 1. GaussianFormer3D models a scene using 3D Gaussians initialized from LiDAR voxel features, updates Gaussians through 3D deformable attention in a LiDAR-camera unified 3D feature space, and finally predicts semantic occupancy via Gaussian-to-voxel splatting. To the best of our knowledge, GaussianFormer3D is the first multi-modal semantic occupancy network that employs a Gaussian-based object-centric scene representation. In summary, our main contributions are as follows:
@@
  • We propose a novel multi-modal Gaussian-based semantic occupancy prediction framework. By integrating LiDAR and camera data, our method significantly outperforms camera-only baselines with similar memory usage.
@@
  • We design a voxel-to-Gaussian initialization module to provide 3D Gaussians with geometry priors from LiDAR data. We also develop an enhanced 3D deformable attention mechanism [31] to update Gaussians by aggregating LiDAR-camera fusion features in a lifted 3D space.
@@
  • We present extensive evaluations on two on-road datasets, nuScenes-SurroundOcc [70] and nuScenes-Occ3D [64], and one off-road dataset, RELLIS3D-WildOcc [83]. Results show that our method performs on par with state-of-the-art dense grid-based methods while having reduced memory consumption and improved efficiency.
2 Related Work
@@
Figure 2: An overview of the proposed GaussianFormer3D framework. We first voxelize LiDAR point clouds to obtain non-empty voxel features for initializing the position and opacity of 3D Gaussians [28]. Then multi-scale LiDAR depth maps and camera feature maps are extracted through projection and an image backbone respectively, and multiplied via outer product to construct a lifted 3D fusion feature space. Gaussians are iteratively updated with 3D sparse convolution, 3D deformable attention, and property refinement. The Gaussian representation is eventually processed by a Gaussian-to-voxel splatting module [25] to generate dense 3D semantic occupancy.
@@
Multi-Modal Semantic Occupancy Prediction. Multi-modal occupancy methods generally outperform single-modal ones since different modalities can complement each other. Common multi-modal sensor configurations for semantic occupancy include LiDAR-camera [52, 66, 32, 3, 63, 60, 87], camera-radar [72, 45] and LiDAR-camera-radar [51]. Among them, LiDAR-camera is the top-performing one as it combines LiDAR’s accurate depth and geometry sensing with camera’s powerful semantic perception. Similar to single-modal pipelines, existing LiDAR-camera occupancy networks mainly build on voxel-based [52, 66, 32, 87, 68] or BEV-based representations [60]. CONet [68] proposes a coarse-to-fine pipeline to sample 3D voxel features for refining the coarse occupancy prediction. Co-Occ [52] obtains multi-modal voxel features through a geometric and semantic-aware fusion module, and employs a NeRF-based implicit volume rendering regularization [49] to enhance the fused representation. OccGen [66] and OccMamba [32] encode multi-modal inputs to produce voxel fusion features, and then decode the features using diffusion denoising [17] and hierarchical Mamba modules [14] respectively to predict semantic occupancy. OccFusion [87] transforms LiDAR and camera inputs into multi-modal voxel features via 2D deformable attention [94] followed by an occupancy head.
@@
3D Gaussians for Autonomous Driving. Due to the inherent advantages of modeling scenes explicitly and continuously, 3D Gaussians [28] have been adopted as the scene representation over the traditional grid-based solutions in 3D semantic occupancy prediction [25, 23, 13] and 4D semantic occupancy forecasting [96]. 3D Gaussians also demonstrated their superiority in real-time image rendering and novel view synthesis and thus have been adopted for driving scene reconstruction and simulation [92, 56, 33, 78, 16]. Furthermore, end-to-end autonomous driving [90] and visual pre-training [74, 85] utilize 3D Gaussians as the driving world representation for various downstream perception and planning tasks. However, these approaches are mainly designed for camera-only autonomous driving, neglecting the potential of multi-modal data in Gaussian initialization and updating. GSPR [55] proposes a Gaussian-based multi-modal place recognition algorithm, and SplatAD [16] designs the first 3D Gaussian splatting pipeline to render both LiDAR and camera data. In this work, we explore utilizing multi-modal data, especially from LiDAR and camera, to learn a better 3D Gaussian representation for more accurate and efficient semantic occupancy prediction.
3 Method
@@
3.1 Scene as 3D Gaussian Representation
Semantic occupancy prediction aims to understand both semantic information and geometric structure of the environment. In the multi-modal scenario, given multi-view camera images =$\mathcal{I}=\{\mathbf{I}_{i}\}_{i=1}^{N_{c}}$ and LiDAR point cloud $\mathcal{P}=\{\mathbf{P}_{i}\}_{i=1}^{N_{p}}$, $\mathbf{P}_{i}=(x_{i},y_{i},z_{i},\eta_{i})$ containing the 3D position and intensity of the point, the goal is to predict the semantic occupancy grid $\mathbf{O}\in\mathcal{C}^{X\times Y\times Z}$, where $N_{c}$, $N_{p}$, $\mathcal{C}$ and $X\times Y\times Z$ represent the number of camera views, the number of LiDAR points, the set of semantic classes and the size of the voxel grid, respectively. Unlike uniform grids in traditional dense grid-based representations, the 3D Gaussian representation can adaptively depict the regions of interest due to the universal approximation capability of Gaussian mixtures [25, 11]. Specifically, a scene is modeled as a set of 3D Gaussians $\mathcal{G}=\{\mathbf{G}_{i}\}_{i=1}^{N_{g}}$, where each Gaussian $\mathbf{G}_{i}$ is parameterized by its mean $\mathbf{m}_{i}\in\mathbb{R}^{3}$, rotation $\mathbf{r}_{i}\in\mathbb{R}^{4}$, scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$, opacity $\sigma_{i}\in[0,1]$ and semantic label $\mathbf{c}_{i}\in\mathbb{R}^{|\mathcal{C}|}$. $N_{g}$ is the total number of Gaussians representing the scene. The value of Gaussian $\mathbf{G}$ evaluated at location $\mathbf{x}$ can be calculated as:

### v2

@@ Abstract
3D semantic occupancy prediction is essential for achieving safe, reliable autonomous driving and robotic navigation. Compared to camera-only perception systems, multi-modal pipelines, especially LiDAR-camera fusion methods, can produce more accurate and fine-grained predictions. Although voxel-based scene representations are widely used for semantic occupancy prediction, 3D Gaussians have emerged as a continuous and significantly more compact alternative. In this work, we propose a multi-modal Gaussian-based semantic occupancy prediction framework utilizing 3D deformable attention, namely GaussianFormer3D. We introduce a voxel-to-Gaussian initialization strategy that provides 3D Gaussians with accurate geometry priors from LiDAR data, and design a LiDAR-guided 3D deformable attention mechanism to refine these Gaussians using LiDAR-camera fusion features in a lifted 3D space. Extensive experiments on real-world on-road and off-road autonomous driving datasets demonstrate that GaussianFormer3D achieves state-of-the-art prediction performance with reduced memory consumption and improved efficiency. Project website: https://lunarlab-gatech.github.io/GaussianFormer3D/.
I Introduction
@@
Perception systems are essential for the development of safe, reliable and intelligent autonomous vehicles and field robots [46]. Among various perception tasks, 3D semantic occupancy prediction is particularly crucial as it enables fine-grained understanding of both geometric and semantic information of the environments [31]. For autonomous driving (AD), recent advances in vision-based occupancy prediction have shown impressive results on large-scale datasets [2, 1]. However, the sensitivity of cameras to lighting variations and their limited depth accuracy still underscore the need to incorporate additional sensing modalities for robust AD.
@@
LiDAR sensors have been widely applied to AD for perception tasks such as 3D object detection [51, 15]. Compared to cameras, LiDAR provides more accurate depth information and finer geometric relationships of objects, making it particularly advantageous for 3D semantic occupancy prediction [30, 5, 40, 39, 43, 24]. However, LiDAR-based pipelines often struggle to capture accurate semantics for small objects, where camera-based methods excel [49]. To balance geometric accuracy and semantic richness, multi-modal fusion algorithms have been proposed to leverage the strengths of complementary sensing modalities. Current approaches include mainly LiDAR-camera fusion [26, 34, 16, 45] and camera–radar fusion [23], with LiDAR-camera fusion showing superior performance and becoming the dominant choice.
@@
Most LiDAR-camera occupancy networks employ a 3D voxel-based [26, 34, 16, 45, 4] representation to model a scene as a dense grid-based structure. Despite achieving comparable performance, these methods inevitably suffer from redundant empty grids and high computational costs. Recently, inspired by the success of 3D Gaussian splatting [13], an object-centric Gaussian representation has been explored in vision-based 3D semantic occupancy prediction, achieving improved computational efficiency. GaussianFormer series [11, 9] represent a scene as a set of 3D Gaussians, each consisting of a mean, covariance and semantics. These Gaussians are refined using a 2D deformable attention [52], and then processed by an efficient Gaussian-to-voxel splatting module to predict semantic occupancy. Despite high efficiency, current Gaussian-based methods [11, 9] rely solely on 2D image to update 3D Gaussians, limiting their ability to model 3D space with accurate depth and fine-grained geometric structure. How to effectively leverage other sensor modalities, such as LiDAR, to refine and obtain a more accurate 3D Gaussian representation for efficient semantic occupancy prediction remains unexplored and challenging.
Figure 1: We propose a new LiDAR-camera fusion-based semantic occupancy prediction framework using 3D Gaussians. We evaluate it on both on-road and off-road driving scenarios. Our method demonstrates superior performance on overall occupancy Intersection-of-Union (IoU), achieves substantial performance gains on small objects (pedestrian, motorcycle) and large surfaces (manmade, vegetation), and consumes less memory during inference.
@@
In this work, we propose GaussianFormer3D: a multi-modal Gaussian-based semantic occupancy prediction framework with 3D deformable attention, as shown in Fig. 1. GaussianFormer3D models a scene using 3D Gaussians initialized from LiDAR voxel features, updates Gaussians through 3D deformable attention in a LiDAR-camera unified 3D feature space, and predicts semantic occupancy via Gaussian-to-voxel splatting. To the best of our knowledge, our model is the first multi-modal semantic occupancy network that employs an object-centric Gaussian-based scene representation. In summary, our main contributions are as follows:
@@
  • We propose a novel multi-modal Gaussian-based semantic occupancy prediction framework. By integrating LiDAR and camera data, ours significantly outperforms camera-only baselines with similar memory usage.
@@
  • We design a voxel-to-Gaussian initialization module to provide 3D Gaussians with geometry priors from LiDAR, and also develop an enhanced 3D deformable attention mechanism to update Gaussians by aggregating LiDAR-camera fusion features in a lifted 3D space.
@@
  • We conduct extensive evaluations on two on-road datasets, nuScenes-SurroundOcc [38] and nuScenes-Occ3D [32], along with an off-road dataset, RELLIS3D-WildOcc [44]. Results show that ours outperforms state-of-the-art dense grid-based methods while achieving reduced memory consumption and improved efficiency.
II Related Work
@@
Figure 2: GaussianFormer3D Overview. We first voxelize LiDAR point clouds to obtain non-empty voxel features for initializing the position and opacity of 3D Gaussians. Then LiDAR depth maps and camera feature maps are extracted respectively, and multiplied via outer product to construct a lifted 3D feature space. Gaussians are iteratively updated with sparse convolution, 3D deformable attention, and property refinement. Gaussians are eventually processed by a Gaussian-to-voxel splatting module to generate semantic occupancy.
@@
Multi-Modal Semantic Occupancy Prediction. Multi-modal occupancy prediction methods generally outperform single-modal approaches, as different modalities provide complementary information. Among them, LiDAR-camera fusion is the top-performing configuration, combining LiDAR’s accurate depth and geometry sensing with the powerful semantic understanding capability of cameras. Similar to single-modal pipelines, most LiDAR-camera occupancy prediction networks are also voxel-based [26, 34, 16, 45, 36]. CONet [36] proposes a coarse-to-fine pipeline to sample 3D voxel features for refining the coarse occupancy prediction. Co-Occ [26] obtains multi-modal voxel features through a geometric and semantic-aware fusion module, and employs a NeRF-based implicit volume rendering regularization [25] to enhance the fused representation. OccGen [34] and OccMamba [16] encode multi-modal inputs to produce voxel fusion features, and then decode the features using diffusion denoising and hierarchical Mamba modules, respectively. OccFusion [45] transforms LiDAR and camera inputs into multi-modal voxel features via 2D deformable attention [52].
@@
3D Gaussians for Autonomous Driving. Due to the inherent advantages of modeling scenes explicitly and continuously, 3D Gaussians [13] have been adopted as the scene representation over the traditional grid-based solutions in 3D semantic occupancy prediction [11, 9, 53]. 3D Gaussians also demonstrated their superiority in real-time image rendering and novel view synthesis, and thus have been adopted for driving scene reconstruction and simulation [50, 42, 8]. Furthermore, end-to-end autonomous driving [48] and visual pre-training [41] utilize 3D Gaussians as the driving world representation for various downstream perception and planning tasks. However, these approaches are mainly designed for camera-only autonomous driving, neglecting the potential of multi-modal data in Gaussian initialization and updating. GSPR [29] proposes a Gaussian-based multi-modal place recognition algorithm, and SplatAD [8] designs the first 3D Gaussian splatting pipeline to render both LiDAR and camera data. In this work, we explore utilizing multi-modal data, especially from LiDAR and camera sensors, to learn a fine-grained 3D Gaussian representation for more accurate and efficient semantic occupancy prediction.
III Method
@@
III-A Scene as 3D Gaussian Representation
Semantic occupancy prediction aims to jointly predict the semantic information and geometric structure of the scene. Given multi-view images $\mathcal{I}=\{\mathbf{I}_{i}\}_{i=1}^{N_{c}}$ and LiDAR point cloud $\mathcal{P}=\{\mathbf{P}_{i}\}_{i=1}^{N_{p}}$, $\mathbf{P}_{i}=(x_{i},y_{i},z_{i},\eta_{i})$ containing the 3D location and intensity of each point, the goal is to predict the semantic occupancy grid $\mathbf{O}\in\mathcal{C}^{X\times Y\times Z}$, where $N_{c}$, $N_{p}$, $\mathcal{C}$ denote the number of camera views, the number of LiDAR points, and the set of semantic classes, and $X\times Y\times Z$ is the size of the voxel grid to be predicted. Unlike uniform grids in traditional grid-based representations, 3D Gaussians can adaptively represent the regions of interest due to the universal approximation capability of Gaussian mixtures [11]. Specifically, a scene is modeled as a set of 3D Gaussians $\mathcal{G}=\{\mathbf{G}_{i}\}_{i=1}^{N_{g}}$, where $N_{g}$ is the total number of Gaussians in a scene. Each Gaussian $\mathbf{G}_{i}$ is parameterized by its mean $\mathbf{m}_{i}\in\mathbb{R}^{3}$, rotation $\mathbf{r}_{i}\in\mathbb{R}^{4}$, scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$, opacity $\sigma_{i}\in[0,1]$ and semantic label $\mathbf{c}_{i}\in\mathbb{R}^{|\mathcal{C}|}$. The value of Gaussian $\mathbf{G}$ evaluated at location $\mathbf{x}$ can be calculated as:
@@ III-A Scene as 3D Gaussian Representation
where $\mathbf{\Sigma}$, $\mathbf{R}$ and $\mathbf{S}$ denote the covariance matrix, rotation matrix and scale matrix. $\rm{diag}(\cdot)$ is the diagonal matrix construction and $\rm{q2r}(\cdot)$ is the quaternion-to-rotation transformation. By summing the contributions of all Gaussians at location $\mathbf{x}$, the occupancy prediction can be formulated as:
The Gaussian-to-voxel splatting module is designed to only aggregate Gaussians within the neighborhood of a targeted voxel instead of querying all Gaussians in a scene to improve efficiency and reduce unnecessary computation and storage [11]. Thus, Eq. 3 can be furthur approximated by replacing $N_{g}$ with $N_{g}(\mathbf{x})$, where $N_{g}(\mathbf{x})$ is the number of neighboring Gaussians at location $\mathbf{x}$. During training, the Gaussian-based occupancy model is trained in an end-to-end manner, supervised by the ground truth semantic occupancy label $\mathbf{\bar{O}}\in\mathcal{C}^{X\times Y\times Z}$. Both cross entropy loss $L_{ce}$ and the lovasz-softmax loss $L_{lov}$ are used for optimization.
III-B Voxel-to-Gaussian Initialization
Two sets of 3D Gaussian features are adopted following GaussianFormer [11]. The first set consists of learnable Gaussian physical properties $\mathcal{G}=\{\mathbf{G}_{i}\in\mathbb{R}^{d}\}_{i=1}^{N_{g}}$ introduced in Sec. III-A, where $d=11+|\mathcal{C}|$, which are also our learning targets. The second set is non-learnable high-dimensional Gaussian features $\mathcal{Q}=\{\mathbf{Q}_{i}\in\mathbb{R}^{m}\}_{i=1}^{N_{g}}$, where $m$ is the feature dimension, serving as queries for the attention mechanism [33] and implicitly encoding the spatial and semantic information during the Gaussian update. Previous work [11] randomly initializes the Gaussian physical properties, and optimizes these properties iteratively through multiple refinement modules. This design constrains Gaussians to learn complex 3D geometry information solely from 2D images, which inevitably encounters inaccurate spatial modeling.
To resolve this issue, we propose a LiDAR-based voxel-to-Gaussian initialization strategy to initialize the mean and opacity of Gaussians with geometry priors from LiDAR data, as indicated in the dashed blue box in Fig. 2. Specifically, we first aggregate the most recent $N_{f}$ LiDAR scans into a combined point cloud $\mathcal{\bar{P}}=\{\mathcal{P}_{i}\}_{i=1}^{N_{f}}$. Then we voxelize the combined point cloud and compute the feature of each non-empty voxel as the mean position and intensity of all points within it. These LiDAR-based voxel features are then used to initialize the position and opacity of 3D Gaussians:
where $i\in\{1,...,N_{g}\}$ denotes the index of Gaussians to be initialized and $v\in\{1,...,N_{v}\}$ denotes the index of all non-empty voxels; $\mathcal{P}_{v}$ is the set of LiDAR points in $\mathcal{\bar{P}}$ falling into voxel $v$. When $N_{g}<N_{v}$, we randomly choose a subset of $N_{g}$ non-empty voxels to initialize Gaussians, otherwise, a subset of $N_{v}$ Gaussians are randomly selected and initialized with non-empty voxels. After initialization, we apply a 3D sparse convolution module to the initialized 3D Gaussians for self-encoding. The features and interactions of Gaussians are efficiently extracted and aggregated through the sparse convolution for updating the Gaussian queries.
III-C LiDAR-Guided 3D Deformable Attention
Lift, Splat, Shoot (LSS) [28] and 2D attention-based methods [52] are widely adopted for feature lifting which transform multi-view 2D images into a 3D space to obtain lifted features. However, LSS suffers from excessive computational costs, hindering its application to multi-scale feature maps that are important for recognizing objects of various sizes. GaussianFormer [11] utilizes a 2D deformable attention (Fig. 3(a)) to extract visual information from 2D images. Despite its efficiency, it suffers from the depth ambiguity problem. As multiple 3D reference points from different Gaussians can be projected to the same 2D position with similar sampling points in the 2D view, this leads to ineffective aggregation, i.e., aggregating the same 2D features for different 3D Gaussian queries. The underlying reason for this is the lack of accurate depth information during the feature lifting and aggregating. A 3D deformable attention operator, namely DFA3D [17], is designed to mitigate the depth ambiguity problem by first expanding 2D feature maps into 3D using estimated depth [6] and then applying an attention mechanism [33] to aggregate features from the expanded 3D feature maps. However, the operator is originally designed for BEV-based 3D object detection (Fig. 3(b)), and relies on DepthNet [6] to estimate monocular depth. Inspired by DFA3D [17], we propose a LiDAR-guided 3D deformable attention mechanism for Gaussian-based semantic occupancy prediction, as illustrated in the dashed orange box in Fig. 2. We first form a unified LiDAR-camera 3D feature space $\mathbf{F}^{{\rm 3D}}$ by conducting outer product between the multi-scale depth maps $\mathbf{F}^{\rm d}$, generated from the LiDAR point cloud, and the multi-scale camera feature maps $\mathbf{F}^{\rm c}:\mathbf{F}^{\rm 3D}=\mathbf{F}^{\rm d}\otimes\mathbf{F}^{\rm c}$. For feature sampling, we design a two-stage key point sampling method (Fig. 3(c)) to aggregate sufficient informative features for updating Gaussian queries. First, we sample a group of 3D reference points $\mathcal{R}_{G}=\{\mathbf{m}_{i}=\mathbf{m}+\Delta\mathbf{m}_{i}|i=1,...,N_{R_{1}}\}$ for each Gaussian $\mathbf{G}$ by shifting its mean $\mathbf{m}$ with learned offsets $\Delta\mathbf{m}$. Then we project these 3D reference points into the fusion feature space $\mathbf{F}^{\rm 3D}$ with extrinsics $\mathcal{T}$ and intrinsics $\mathcal{K}$, where each projected reference point is positioned at $\mathbf{\bar{m}}_{i}=(u_{i},v_{i},d_{i})$.
Figure 3: Comparison of different feature sampling methods.
After projection, we further generate learnable sampling offsets $\Delta\mathbf{\bar{m}}_{ij}=(\Delta u_{ij},\Delta v_{ij},\Delta d_{ij})$ for each projected reference point $\mathbf{\bar{m}}_{i}$. The overall sampling points of a given Gaussian $\mathbf{G}$ in the fusion feature space $\mathbf{F}^{\rm 3D}$ can be formulated as: